FEL Upgrade Project Weekly Brief
December 9-13, 2002


We have survived another extremely busy and productive week of installation and commissioning activities. Highlights include: good progress on the injector RF system; magnet control pages in the front end of the machine are being successfully brought to life; the gun bakeout was started (but then had to be interrupted as we opened a leak which was quickly identified in a feedthrough on the downstream beamline); the first reverse bend (GQ) magnet successfully met our demanding specifications ; all vacuum chambers for both recirculation arcs are now on-site; the remaining magnets for the UV FEL (6 optical chicane, GW magnets) arrived this week; and we had a successful test of our cryo-cooled mirror concept last weekend for our JTO project.

We had a successful quarterly ES&H inspection this week. The results were reviewed by the project team with action items on several minor items distributed to responsible parties. A response to several questions from ONR related to FY03 funding was prepared and sent to our DOE contract monitor today. The next increment in FY03 funding ($761k) was sent on Dec.12th from ONR. We thank ONR for their timely handling of these budget matters.

On Monday, Dec. 11th, a conference call was held between JLab and Aerospace Corporation to finalize contractual matters on the existing and planned joint statements of work for our AFRL funded collaboration on UV FEL and processing tool development.

On Tuesday, Dec 10th, George Neil and Michelle Shinn gave presentations to the JTO Program Office in Albuquerque on the close-out of our Phase 1 (FY02) projects on : (1) advanced drive laser development and (2) cryo-cooled high power mirrors. The review was successful and the JTO program office has forwarded a recommendation to the Navy JTO contract monitor that the Phase 2 (FY03) JTO funds for these projects be released.

On Thursday, Dec. 12th, Jefferson Lab hosted the quarterly meeting of Virginia Research and Technology Advisory Commission (VRTAC) which included a tour of the FEL. VRTAC has helped raise the state-wide visibility of the FEL program in the past. Virginia’s Secretary of Technology, George Newstrom, and Deputy Secretary of Education, Peter Blake, were in attendance at the meeting.

WBS 4 (Injector):
Set up and started a bake on the gun and downstream beam line. Pumping was being done into a turbo pump to help remove residual hydro carbon contamination from the system before we move over to the internal ion pumps. While at 150C a leak opened up on a feedthrough on the BPM/ bellows assembly. The decision was made to weld the feedthrough over for the moment and build another offline for future replacement.

Gun HVPS - Work continues with the I&C group on the customer interface control box. Good progress is being made.

WBS 6 (RF):
Quarter Klystrons - Unit #4 klystron and HVPS were operated to 40 kW into shorts with an external RF source. The PSS by-pass jumpers were removed and the Quarter re-certified under PSS control. The final wiring was completed for operation with the RF Control Modules. It will be tested today at noon. The next step is SRF qualification.

WBS 8 (Instrumentation):

Fast and productive! Another week has flown by. The many details are being nailed down one after another. All of the scanner hardware was moved to the injector area awaiting the completion of the bake. The rail and final adjustments will be made once the surrogate (GreenNe) laser is mounted. The image shows the scan done with a FET/Load resistor. The image is rotated because the old light box port was at 45 degrees. The colors represent variations in Qe. Each scan takes 64k data points, 16 bits deep of X axis, Y axis, and current response in 32 seconds.

Final connections for the trim magnets and injection GV trim coils were completed in the 0F region except for the air cores magnets up stream of the 1/4 cryo. These magnets had to be disconnected and unassembled to support the gun bake. Final high level checkout of the 0F region will begin next week. Work has begun on the 3F and 4F trim magnet software, 1st Arc shunts modules and respective Ops screens.

First cut of the Magnet system block diagram was received from EECAD and reviewed. Commissioning of the HV interlock chassis has begun. The digital and analog HV I/O module artwork has been received from EECAD and under review. Test cables were pulled in for the High Reflector LVDT's and terminated.

Beam Loss Monitors (BLMs) - The prototype and the production card are tested and installed into iocfel8. The accompanying break-out boards are installed as well in the back of the crate. As each board pair provides 12 channels of the system, we have 24 channels ready to be connected to the blm heads and we will proceed with the "in-system" testing when time permits. The high-level software is still under development and is now our main area of development for this system.

Drive Laser Controls and Pulse Controller Upgrade (DLPC) - The VME based micropulse generator board is installed into iocfel10 and is being tested "in-system". All the bench tests have been successful. The vme crate has been fitted with -5V power supply to support the board. A new scope has been installed into FL06B01 in the clean room. This is a dedicated scope for the drive laser's and the cathode's photo-diode signals. See related images at http://laser.jlab.org/wbs8. A MPS summation chassis design was completed and fabrication has begun. This will combine all of the MPS signals from the control room and port them to the MPS rack over a single cable. This will also make troubleshooting easier. Until the MPS is fully certified, the DLPC's MPS inputs have been hardwired to Beam Mode 2. This will limit running to 2Hz, 250us line (AC)-locked, pulsed mode. This is the maximum safe operating limit until the recirculation path is complete in early spring. This will allow for complete commissioning of the wiggler and optical cavity.

The dump water skid conductivity/temperature monitoring chassis is working and setpoints for alarms and the polishing loop have been programmed. Drawing revisions are in progress. The first Synchrotron Light Monitor (SLM) attenuator is complete, the additional 9 will now be fabricated.

Minor support was given to the magnet test facility, there were a couple of connector problems. They still have no real electrical support.

The quarterly EH&S safety inspection was conducted on Wednesday. Overall we looked very good. Individuals with items requiring attention will be notified by the safety warden.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
• The DY magnet may be assembled later today as the qualified personnel become available during the Gun’s bake-out.
• The stands for the DY are installed and aligned.
Diversion of the beam after the Wiggler.
• The vacuum group is working on the vacuum pipes and the stand in the extraction area beyond the wiggler.
• The GW magnet, planned to go into the extraction position is in the enclosure and ready to be placed on its stand.
Optical Chicane Dipoles (GW)
• The 6 GW Dipoles for the UV machine were received. This completes the UV magnet inventory.
Injector Dipoles (DU/DV)
• No action
Arc 180 Degree Dipoles (GY)
• Wang NMR completed the impregnation of the first coil and cured it. It is being extracted from the mold and the
    second coil is being prepared,
•  Wang NMR is glued the Purcell gap shims on top yoke.
•  We started mounting the third power supply in the Magnet Test Area so that they can test three different magnets in
    parallel. This gets them through the crunch coming up where a quadruple from Physics as well as the CEBAF
    recirculation experiment dipoles as well as the FEL’s GY all require testing at once.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
• The first GQ obtained David Douglas’ ± 1 part in 10,000 criteria for field integral as well as core field at 80 and 145
    MeV/c. However, the pass at the low momentum could only be accomplished for core field by using a different
    (120 A) hysteresis loop below 115 MeV/c operation. David prefers simpler operation. The alternative is trim coils on
    the pole face with their added complexity or possible correction by the installed sextupoles.
• We placed the first GX on the test stand instead of the GQ in order to test out if it requires the same hysteresis loops as
    the GQ in order to meet David’s criteria. The first tests were run and the data is being evaluated.
• Wang is ready to impregnated the fourth GX coil.
Trim Quad (QT)
• The trim coil to create the GC vertical corrector function within the trim quad is out for bid.
Sextupole (SF)
•  No progress
Octupole (OT)
•  We signed off the Specification and have started the procurement paperwork but will hold procurement pending funds
    on this descoped item.
Beam Line and Vacuum
•  All the remaining chambers for the Arcs were received from Master Machine.
•  Optical Chicane chamber fabrication in the JLab Shop continued
•  We broke vacuum briefly in the injection area to insert the vacuum chamber mimic units in the first two GV dipoles
    and to adjust the X Chamber to a position that creates more aperture.

WBS 11 (Optics):
Work this week continued on establishing the utilities (electrical and cooling) to the optical cavity internals. The lower limit switches (UHV-rated) were wired and set in their proper position on the foundation plates. The leadscrews and leadscrew nuts were changed out as well, since all of our testing in the summer had produced a moderate amount of wear. These activities require Survey and Alignment to realign the foundation plates in the vacuum vessels, and this work begins today (12/13). The flexible tubes that the cooling water flows through were mounted on their conveyor chains. We plan on mounting them in the vacuum vessels next week, or the first of the year. The internal wiring harnesses for the LVDTs for the HR are complete and ready for installation. Fabrication of the diagnostic viewer/laser safety shutter and the insertable mirrors are being coordinated by our Machine Shop. They still plan to deliver parts before Christmas, and already we are receiving some of those parts.

Our cryocooled mirror tests continued into the weekend, and were quite successful. We presented those results to the JTO on 12/10. Today we host visitors from DARPA and a laser vendor.